# Kernel-U-Net: Hierarchical and Symmetrical Framework for Multivariate Time Series Forecasting

### 03 Jan 2024 v1

@@ Abstract
Time series forecasting task predicts future trends based on historical information. Recent U-Net-based methods have demonstrated superior performance in predicting real-world datasets. However, the performance of these models is lower than patch-based models or linear models. In this work, we propose a symmetric and hierarchical framework, Kernel-U-Net, which cuts the input sequence into slices at each layer of the network and then computes them using kernels. Furthermore, it generalizes the concept of convolutional kernels in classic U-Net to accept custom kernels that follow the same design pattern. Compared to the existing linear or transformer-based solution, our model contains 3 advantages: 1) A small number of parameters: the parameters size is $O(log(L)^{2})$ where $L$ is the look-back window size, 2) Flexibility: its kernels can be customized and fitted to the datasets, 3) Computation efficiency: the computation complexity of transformer modules is reduced to $O(log(L)^{2})$ if they are placed close to the latent vector. Kernel-U-Net accuracy was greater than or equal to the state-of-the-art model on six (out of seven) real-world datasets.
@@ 1 Introduction
Over the past few decades, time series forecasting solutions have evolved from traditional statistical methods (e.g., ARIMA Khandelwal et al. (2015) and machine learning techniques (e.g., GBRT Persson et al. (2017)) to deep learning-based solutions, such as RNN recurrent neural networks Tokgöz and Ünal (2018), LSTM Kong et al. (2017), Temporal Convolutional Network (TCN) Hewage et al. (2020) and Transformer-based model Li et al. (2019).
Large-scale time series forecasting addresses the problem of predicting a longer future horizon with a larger look-back window. With the increasing size of data, the cost of learning deep models increases crucially. Moreover, models such as LSTM or RNN suffer gradient vanishing problems in learning long sequences. To balance the quality of prediction and efficiency of learning deep models, researchers proposed hierarchical structure in Duan et al. (2022), pyramidal structure in Pyraformer Liu et al. (2022a), AutoEncoders Asres et al. (2022), U-Net-like model in Yformer Madhusudhanan et al. (2023) or patches in PatchTST Nie et al. (2023). Meanwhile, there is an increasing trend in combining different components for learning particular time series. For example in time series classification, researchers studied hybrid models such as combing GRU Elsayed et al. (2019), LSTM Lin et al. (2020); Khan et al. (2021) or Transformer Huang et al. (2020b) with linear or CNN layers Hewage et al. (2020).
@@
Among the Transformer models applying to time series data, Informer Zhou et al. (2021), Autoformer Wu et al. (2021), and FEDformer Zhou et al. (2022) are the best variants that improve the quality of prediction. Meanwhile, a recent paper Zeng et al. (2023) gives a simple linear model NLinear which outperforms pure transformer-based models on a variety of common benchmarks and challenges the Transformer models for time series forecasting. Later, the authors in PatchTST Nie et al. (2023) argued that applying transformer modules on a point-wise sequence degrades the performance. By adding a linear patch layer, they successfully relieved the overfitting problem of transformer modules and they reached state-of-the-art result on several large datasets.
We remark that models are particularly sensible to different types of datasets. For example, NLinear is more robust against overfitting in small-size datasets such as Etth and EttmZeng et al. (2023) and PatchTST is more expressive in large-size datasets such as Weather and Traffic etc.Nie et al. (2023). At the same time, it is also important to notice the parameter size in Linear models such as NLinear is in $O(L^{2})$ and the computation complexity is $O(L)$. Meanwhile, the parameter size in a transformer block is $O(1)$ but the computational complexity is $O(L^{2})$ in general Transformer modules. It is reduced to $O(Llog(L))$ using Propsparse and to $O(\frac{L}{S}log(\frac{L}{S}))$ using patches, where $S$ is the size of the patch. Based on previous analysis, there is a motivation to propose a more flexible framework for combining the advantages of different modules in high-paced development and testing.
As a classical and highly expressive model in medical image segmentation, Convolutional U-net has an elegant symmetrical encoder and decoder structureRonneberger et al. (2015). It coincides well with the time series forecasting problem because the inputs and outputs are drawn from the same distribution. The first U-Net-like model for time series forecasting was YformerMadhusudhanan et al. (2023), which used Transformers in the encoder and decoder. As mentioned before, applying a Transformer on point-wise data can lead to overfitting. Therefore, we are interested in answering the question of whether there is a U-net-like structure that performs well for the time series forecasting task.
Figure 1: Structure of Kernel U-Net
We propose a general hierarchical and symmetrical framework, Kernel-U-Net, inspired by convolutional U-net Ronneberger et al. (2015) and Yformer Madhusudhanan et al. (2023) for time series forecasting. Kernel-U-Net offers $O(log(L)^{2})$ parameters size for processing a time series of size $L$. Meanwhile, it provides convenience for composing particular models with custom kernels that follow the same design pattern. Since the hierarchical structure reduces the length of the sequence at each level exponentially, the complexity of learning such sequences is simultaneously reduced. In particular, the computation complexity of Transformer and Long Short-term Memory(LSTM) modules is reduced to $O(log(L)^{2})$ and $O(log(L))$ if they are only employed in the layers close to the latent vector. Therefore, Kernel-U-Net helps reduce the computation cost of the Transformer and relieves the gradient vanishing problem in LSTM in learning long sequences.
@@
To fully study the performance and efficiency of Kernel-U-Net, we conduct experiments for time series forecasting tasks on several widely used benchmark datasets. Our results show that in time series forecasting, Kernel-U-Net meets or outperforms baseline models, such as NLinear and PatchTST, in the majority cases.
@@
  • We propose Kernel-U-Net, a general hierarchical and symmetrical framework that partitions input sequences into small slices and learns them at each layer of the network.
  • Kernel-U-Net generalizes the concept of kernels and provides convenience for composing particular models with custom kernels that follow the same design pattern. The parameter size is guaranteed in $O(log(L)^{2})$.
  • The computation complexity of Transformer module is $O(Llog(L))$ and can be reduced to $O(log(L)^{2})$ if it is only employed in the layers close to the latent vector.
  • Kernel-U-Net achieves results with MSE close to or better than the state-of-the-art methods on 6 (out of 7) time series data sets.
We conclude that Kernel-U-Net is a highly potential candidate for large-scale time series forecasting tasks. It is a general hierarchical and symmetrical framework with low complexity in terms of parameters. It reaches state-of-the-art results in the majority cases. With flexible customized kernels, it adapts well in high-paced development and test context.
@@ 2 Related works
Transformer Vaswani et al. (2017) was initially introduced in the field of Natural Language Processing(NLP) on language translation tasks. It contains a layer of positional encoding, blocks that are composed of layers of multiple head attentions, and a linear layer with SoftMax activation. As it demonstrated outstanding performance on NLP tasks, many researchers follow this technique route. Bidirectional Encoder Representations from Transformers (BERT)Devlin et al. (2019) is designed to pre-train deep bidirectional representations from the unlabeled text by joint conditioning on both left and right context in all layers. Generative Pre-trained Transformer 3(GPT-3) pre-trained a massive transformer-based model using 175 billion parameters Brown et al. (2020) and achieved strong performance on different types of downstream natural language tasks without requiring any fine-tuning.
Later, Vision Transformers (ViTs) Dosovitskiy et al. (2021) ViT applies a pure transformer directly to sequences of image patches to classify the full image and outperformed CNN based method on ImageNetDeng et al. (2009). Swin Transformer Liu et al. (2021) proposed a hierarchical Transformer whose representation is computed with shifted windows. As a shifted window brings greater efficiency by limiting self-attention computation to non-overlapping local windows, it also allows cross-window connection. This hierarchical architecture has the flexibility to model at various scales and has linear computational complexity concerning image size.
In time series forecasting, the researchers were also attracted by transformer-based models. LogTrans Li et al. (2019) proposed convolutional self-attention by employing causal convolutions to produce queries and keys in the self-attention layer. It proposes also a LogSparse Transformer with only $O(L(logL)^{2})$ space complexity to break the memory bottleneck. Informer Zhou et al. (2021) has an encoder-decoder architecture that validates the Transformer-like model’s potential value to capture individual long-range dependency between long sequences. It proposes a ProbSparse self-attention mechanism to efficiently replace the canonical self-attention. It achieves the $O(LlogL)$ time complexity and $O(LlogL)$ memory usage on dependency alignments. PyraformerLiu et al. (2022a) simultaneously captures temporal dependencies of different ranges in a compact multi-resolution fashion. Theoretically, by choosing parameters appropriately, it achieves concurrently the maximum path length of $O(1)$ and the time and space complexity of $O(L)$ in forward pass.
@@
Recently, YFormer Madhusudhanan et al. (2023) proposed a U-Net-like architecture that combines coarse and fine-grained features to capture long-range effects on all scale levels. PatchTST Nie et al. (2023) adopted the patches method from ViTs and proposed a time series transformer with a linear patch layer. It reaches state-of-the-art results on many time series forecasting benchmarks. The author claims Reduction in time and space complexity of learning transformer block from $O(L^{2})$ to $O((\frac{L}{S})^{2})$, where $L$ is sequence size and $S$ is stride size.
@@
U-Net is a neural network architecture designed primarily for medical image segmentation Ronneberger et al. (2015). U-net is composed of an encoder and a decoder. At the encoder phase, a long sequence is reshaped into small slices and reduced by kernels step by step into a latent vector. At the decoder stage, the vector is developed by kernels for generating an output with the same shape as the input. With the help of skip-connection between the encoder and decoder, U-Net can also capture and merge low-level and high-level information easily.
With such a neat structural design, U-Net has achieved great success in a variety of applications such as medical image segmentation Ronneberger et al. (2015), biomedical 3D-Image segmentation Çiçek et al. (2016), time series segmentation Perslev et al. (2019), image super-resolution Han et al. (2022) and image denoising Zhang et al. (2023). Its techniques evolve from basic 2D-U-Net to 3D U-Net Çiçek et al. (2016), 1D-U-Net Azar et al. (2022), Weighted Res-UNet Xiao et al. (2018) with residual structure and weighted attention mechanism, U-Net++ Zhou et al. (2018) with nested U-net inside each layer, UNet3+ Huang et al. (2020a) that allows full-scale skip connections incorporating low-level details with high-level semantics from feature maps in different scales and Swin-U-Net Cao et al. (2023) that only use transformer blocks at each layer of U-Net. Modification on U-Net layers has been studied by Selective Kernel U-Net Byra et al. (2020) which merges information from different sizes of convolutional kernels and Large Kernel U-Net Jia et al. (2022) which combines TransMorph layers for better performance in medical image segmentation.
@@
In time series processing, learning sequences of length $L$ requires $O(L)$ computations for recurrent models such as RNN, LSTM, GRU, and $O(L^{2})$ computations for transformer-based models. With ProbSparse Attention, the complexity can be reduced to $O(Llog(L))$. To balance the expressivity of complex modules and computational efficiency, researchers investigated hierarchical architecture and hybrid methods that merge different modules into the network.
For example, authors in Du et al. (2015) investigated a tree structure model made of bidirectional RNN layers and concatenation layers for skeleton-based action recognition, authors in Xiao and Liu (2016) stacked RNN and LSTM with Attention mechanism for semantic relation classification of text, authors in Kowsari et al. (2017) applied hierarchical LSTM and GRU was used for document classification. In time series processing, the authors in Hong and Yoon (2017) combine Deep Belief Network (DBN) and LSTM for sleep signal pattern classification, authors in LogTrans Li et al. (2019), Informer Zhou et al. (2021)) and Pyraformer Liu et al. (2022a) studied hierarchical or pyramidal transformers for time series forecasting.
Kernel-U-Net is a general hierarchical and symmetrical framework with low complexity in terms of parameter size. By replacing the convolutional kernel in classical U-Net, our model allows users to define custom kernels such as Attention, LSTM, or simple linear multiplication with a non-linear activation function for extracting features on complex time series. There are 3 advantages in Kernel-U-Net: 1) A small number of parameters: the total parameter of model is in order of $O(log(L)^{2})$, 2) Flexibility: the kernels can be replaced easily by arbitrary function that follows the same design mode so that non-linear kernels can be easily combined into it, 3) Computation efficiency: the computation complexity of transformer modules is reduce to $O(log(L)^{2})$ if they placed close to latent vector.
Algorithm 1 Kernel Wrapper
Input: kernel $f_{i}$, input segments $x$, input length $L_{in}$ and dimension $M_{in}$, output length $L_{out}$ , and dimension $M_{out}$
Output: Instance of Kernel Wrapper
1:  # Initiation of main kernel function $f$ is
2:  $f=$ $f_{i}(L_{in},M_{in},L_{out},M_{out})$
3:  # Holds variables for global operation
4:  _unet_skip_output = None
5:  _unet_skip_input = None
6:
7:  # Define the forward function as :
8:  Reshape $x$ into (-1, $L_{in}$,$M_{in}$)
9:  if _unet_skip_input is not None then
10:     $x=x+\_unet\_skip\_input$
11:  end if
12:  # Apply $f$ on $x$
13:  $y=f_{i}(x)$
14:  _unet_skip_output = $y$
15:  return $y$
@@ 3.1 Problem Formulation
Time series forecasting.
In the case of time series forecasting, the dataset is composed of a time series of characteristics $x$ and future series $\hat{x}$. Let $x_{t}$ be the feature at the time step $t$ and $L$ the length of the look-back window. Given a historical data series $(x_{t+1},...,x_{t+L})$ of length $L$, time series forecasting task is to predict the value $(\hat{x}_{t+L+1},...,\hat{x}_{t+L+T})$ in the future T time steps. Then we can define the basic time series forecasting problem: $(\hat{x}_{t+L+1},...,\hat{x}_{t+L+T})=f(x_{t+1},...,x_{t+L})$ , where $f$ is the function that predicts the value $(\hat{x}_{t+L+1},...,\hat{x}_{t+L+T})$ in the future T time steps based on a series $(x_{t+1},...,x_{t+L})$.
Algorithm 2 Kernel U-Net Encoder
Output: Instance of Kernel U-Net Encoder
1:  $L=[L_{1},L_{2},...,L_{n},$ $L_{n+2},...,L_{n+m}]$
2:  # Create first layer
[…]
4:  # Create intermediate layers
[…]
6:     layers.append(KernelWrapper($f_{i},L_{i},M_{i},1,M_{i+1}$))
7:  end for
8:  # Create last layer
9:  layers.append(KernelWrapper($f_{n+m},L_{n+m},M_{n+m},$ $L_{h},M_{h}$))
10:  encoder = nn.Sequential (*layers)
@@ 3.2 Kernel U-Net
Kernel U-Net follows the hierarchical structure of convolutional U-Net. It accepts a custom kernel to replace the linear kernel inside the original convolutional U-Net. More precisely, it repeatedly applies custom kernels on small slices of large-scale input data. In this context, Kernel U-Net avoids problems in processing long sequences such as gradient vanishing in recurrent models and quadradic complexity in transformers-based models. In the following paragraphs, we describe methods such as the hierarchical partition of the input sequence, and the creation of kernel wrapper, encoder, decoder, and U-net.
@@
Let us consider an input trajectory matrix $X_{t}\in R^{L\times M}$, where $L$ is lookback window size, $M$ is feature size and $t$ is the time step. Given a list of multiples $(L_{1},...,L_{n})$ for look-back window and $(M_{1},...,M_{m})$ for feature such that $L=\prod_{1}^{n}L_{i}$ and $M=\prod_{1}^{m}M_{i}$ We reshape $X_{t}$ as small slices $X_{t,i,j}\in R^{L_{1}\times M_{1}}$ such that the concatenation of all slices is equal to its original trajectory matrix:
@@
The partitioned sequences are processed by kernels in the encoder. Since the operations in the decoder are symmetrical to those in the encoder, there will also be output $\hat{X}_{t+L}$ composed of generated segments $\hat{X}_{t+L,i,j}$.
@@
At the stage of the encoder, the shape of input $X$ is $(B,L,M)$. Let us assume that $M_{hidden}$ is the unique dimension of the hidden vector at each intermediate layer to simplify the problem. We reshape $X$ into $(B,\prod_{2}^{n}L_{i},L_{1},\prod_{2}^{m}M_{i},M_{1})$ then transpose it into $(B,\prod_{2}^{m}M_{i},\prod_{2}^{n}L_{i},L_{1},M_{1})$ and reshape it into $(B*\prod_{2}^{m}M_{i}*\prod_{2}^{n}L_{i},L_{1},M_{1})$ The kernel $f_{L_{1}}$ can now process the reshaped $X$ as a large batch of slices and output a vector $Z_{L_{1}}$ in shape of $(B*\prod_{2}^{m}*M_{i}*\prod_{2}^{n}L_{i},1,M_{hidden})$. After this operation, we reshape the output $Z_{L_{1}}$ into $(B*\prod_{2}^{m}M_{i}*\prod_{3}^{n}L_{i},L_{2},M_{hidden})$ for next layer. Iteratively, the encoder of Kernel U-Net processes all the multiples in the list $(M_{m},...,M_{2},L_{n},...,L_{2})$ in reverse order, and gives the final latent vector $Z$ in the shape of $(B,1,M_{hidden})$ as a batch of the latent vector.
At the stage of the decoder, the operations are reversed. We send a latent vector Z of shape $(B,1,M_{hidden})$ to decoder $f_{M_{m}}$ and get an output $Z_{M_{m}}$ in shape of $(B,M_{m},M_{hidden})$, then we reshape it into $(B*M_{m},1,M_{hidden})$ for the next kernels. At the end of this iteration over multiples in the list $(M_{m-1},...,M_{2},L_{n},...,L_{2},L_{1})$, we have finally a vector in the shape of $(B*\prod_{2}^{m}M_{i}*\prod_{2}^{n}L_{i},L_{1},M_{1})$. The last operations are reshaping it into $(B,\prod_{2}^{m}M_{i},\prod_{2}^{n}L_{i},L_{1},M_{1})$ , transposing it into the vector in shape $(B,\prod_{2}^{n}L_{i},L_{1},\prod_{2}^{m}M_{i},M_{1})$ and reshaping it to shape $(B,L,M)$.

### 12 Feb 2024 v2

# Kernel-U-Net: Symmetric and Hierarchical Architecture for Multivariate Time Series Forecasting

@@ Abstract
Time series forecasting task predicts future trends based on historical information. Transformer-based U-Net architectures, despite their success in medical image segmentation, have limitations in both expressiveness and computation efficiency in time series forecasting as evidenced in YFormer. To tackle these challenges, we introduce Kernel-U-Net, a symmetric and hierarchical U-shape neural network architecture. The kernel-U-Net encoder compresses gradually input series into latent vectors, and its symmetric decoder subsequently expands these vectors into output series. Specifically, Kernel-U-Net separates the procedure of partitioning input time series into patches from kernel manipulation, thereby providing the convenience of executing customized kernels. Our method offers two primary advantages: 1) Flexibility in kernel customization to adapt to specific datasets; 2) Enhanced computational efficiency, with the complexity of the Transformer layer reduced to linear. Experiments on seven real-world datasets, considering both multivariate and univariate settings, demonstrate that Kernel-U-Net’s performance either exceeds or meets that of the existing state-of-the-art model PatchTST in the majority of cases and outperforms Yformer. The source code for Kernel-U-Net will be made publicly available for further research and application.
@@ 1 Introduction
Over the past few decades, time series forecasting solutions have evolved from traditional statistical methodsKhandelwal et al. (2015) and machine learning techniquesPersson et al. (2017) to deep learning-based solutions, such as recurrent neural networks(RNN) Tokgöz and Ünal (2018), Long Short-term Memory (LSTM) Kong et al. (2017), Temporal Convolutional Network (TCN) Hewage et al. (2020) and Transformer-based model Li et al. (2019).
@@
Among the Transformer models applying to time series data, Informer Zhou et al. (2021), Autoformer Wu et al. (2021), and FEDformer Zhou et al. (2022) are the best variants that incrementally improved the quality of prediction. As a recent paper Zeng et al. (2023) challenges the efficiency of Transformer-based models with a simple linear layer model NLinear, the authors in Nie et al. (2023) argued that the degrades of performance comes from the wrong application of transformer modules on a point-wise sequence and the ignorance of patches. By adding a linear patch layer, their model PatchTST successfully relieved the overfitting problem of transformer modules and reached state-of-the-art results.
We observe that models display distinct strengths depending on the dataset type. For instance, NLinear stands out for its efficiency in handling univariate time series tasks, particularly with small-size datasets. On the other hand, PatchTST is noteworthy for its expressiveness in multivariate time series tasks on large-size datasets. These contrasting attributes highlight the necessity for a unified but flexible architecture. This architecture would aim to integrate various modules easily, allowing for specific customized solutions. Such integration should not only ensure a balance between computational efficiency and expressiveness but also respond to requests for rapid development and testing.
The Convolutional U-net, as a classic and highly expressive model in medical image segmentationRonneberger et al. (2015), features a symmetric encoder and decoder structure that is elegant in its design. This model’s structure is particularly suited to the time series forecasting task, as both inputs and outputs in this context are typically derived from the same distribution. The first U-shape model adapted for time series forecasting was the YformerMadhusudhanan et al. (2023), which incorporated Transformers in both its encoder and decoder components. As mentioned previously, employing Transformers on point-wise data has the potential to cause overfitting issues. Therefore, our investigation aims to discover if there is a U-shape architecture effective in time series forecasting, which also possesses the capability to integrate various modules flexibly, thus facilitating the specific customization of solutions.
To tackle this problem, we propose a symmetric and hierarchical architecture, Kernel-U-Net (K-U-Net), inspired by convolutional U-net and Yformer for time series forecasting. K-U-Net provides convenience for composing particular models with custom kernels that follow the design pattern. By replacing linear kernels with transformer or LSTM kernels, the model gains enhanced expressiveness, allowing it to capture more complex patterns and dependencies in the data. Furthermore, the hierarchical structure of K-U-Net exponentially reduces the input length at each level, thereby concurrently decreasing the complexity involved in learning such sequences. Notably, when Transformer modules are utilized in the second or higher-level layers, the computation cost remains linear, ensuring efficiency in processing.
@@
To fully study the performance and efficiency of K-U-Net, we conduct experiments for time series forecasting tasks on several widely used benchmark datasets. We compose 30 variants of K-U-Net by placing different kernels at different layers and then we choose the best candidates for fine-tuning. Our results show that in time series forecasting, K-U-Net exceeds or meets the state-of-the-art results, such as NLinearZeng et al. (2023) and PatchTST, in the majority of cases.
@@
  • We propose Kernel-U-Net, a symmetric and hierarchical architecture that partitions input sequences into small patches at each layer of the network.
  • Kernel-U-Net generalizes the concept of kernels and provides convenience for composing particular models with custom kernels that follow the design pattern.
  • The computation complexity is guaranteed in linear when employing Transformer kernels at the second or higher layers.
  • Kernel-U-Net exceeds or meets the state-of-the-art results in most cases.
We conclude that Kernel-U-Net stands out as a highly promising option for large-scale time series forecasting tasks. Its symmetric and hierarchical design provides a balance of low computational complexity and high expressiveness. In most scenarios, it either surpasses or is slightly below the state-of-the-art results. Furthermore, its adaptability in fast-paced development and testing environments is ensured by the use of flexible, customizable kernels.
Figure 1: Comparison of the structures of different U-Net. a) classic Convolutional U-Net, b) Kernel U-Net, c) Swin Transformer U-Net
@@ 2 Related works
Transformer Vaswani et al. (2017) was initially introduced in the field of Natural Language Processing(NLP) on language translation tasks. It contains a layer of positional encoding, blocks that are composed of layers of multiple head attentions, and a linear layer with SoftMax activation. As it demonstrated outstanding performance on NLP tasks, many researchers follow this technique route.
Vision Transformers (ViTs) Dosovitskiy et al. (2021) ViT applies a pure transformer directly to sequences of image patches to classify the full image and outperformed CNN based method on ImageNetDeng et al. (2009). Swin Transformer Liu et al. (2021) proposed a hierarchical Transformer whose representation is computed with shifted windows. As a shifted window brings greater efficiency by limiting self-attention computation to non-overlapping local windows, it also allows cross-window connection. This hierarchical architecture has the flexibility to model at various scales and has linear computational complexity concerning image size.
In time series forecasting, the researchers were also attracted by transformer-based models. LogTrans Li et al. (2019) proposed convolutional self-attention by employing causal convolutions to produce queries and keys in the self-attention layer. To improve the computation efficiency, the authors propose also a LogSparse Transformer with only $O(L(logL)^{2})$ space complexity to break the memory bottleneck. Informer Zhou et al. (2021) has an encoder-decoder architecture that validates the Transformer-like model’s potential value to capture individual long-range dependency between long sequences. The authors propose a ProbSparse self-attention mechanism to replace the canonical self-attention efficiently. It achieves the $O(LlogL)$ time complexity and $O(LlogL)$ memory usage on dependency alignments. PyraformerLiu et al. (2022a) simultaneously captures temporal dependencies of different ranges in a compact multi-resolution fashion. Theoretically, by choosing parameters appropriately, it achieves concurrently the maximum path length of $O(1)$ and the time and space complexity of $O(L)$ in forward pass. PatchTST employed PropSparse and patches, thus reducing computation complexity to $O(\frac{L}{S}log(\frac{L}{S}))$, where $S$ is the size of the patch.
@@
U-Net is a neural network architecture designed primarily for medical image segmentation Ronneberger et al. (2015). U-net is composed of an encoder and a decoder. At the encoder phase, a long sequence is reduced by a convolutional layer and a max-pooling layer gradually into a latent vector. At the decoder stage, the latent vector is developed by a transposed convolutional layer for generating an output with the same shape as the input. With the help of skip-connection between the encoder and decoder, U-Net can also capture and merge low-level and high-level information easily.
With such a neat structural design, U-Net has achieved great success in a variety of applications such as medical image segmentation Ronneberger et al. (2015), biomedical 3D-Image segmentation Çiçek et al. (2016), time series segmentation Perslev et al. (2019), image super-resolution Han et al. (2022) and image denoising Zhang et al. (2023). Its techniques evolve from basic 2D-U-Net to 3D U-Net Çiçek et al. (2016), 1D-U-Net Azar et al. (2022) and Swin-U-Net Cao et al. (2023) that only use transformer blocks at each layer of U-Net.
@@
In time series processing, the increasing size of data degrades the performance of deep models and increases crucially the cost of learning them. For example, recurrent models such as RNN, LSTM, GRU have a linear complexity but suffer the gradient vanishing problem when input length increases. Transformer-based block captures better long dependencies but requires $O(L^{2})$ computations in general. To balance the expressiveness of complex models and computational efficiency, researchers investigated hybrid models that merge different modules into the network and hierarchical architectures.
For example, authors in Du et al. (2015) investigated a tree structure model made of bidirectional RNN layers and concatenation layers for skeleton-based action recognition, authors in Xiao and Liu (2016) stacked RNN and LSTM with Attention mechanism for semantic relation classification of text, authors in Kowsari et al. (2017) applied hierarchical LSTM and GRU for document classification. In time series processing, the authors in Hong and Yoon (2017) combine Deep Belief Network (DBN) and LSTM for sleep signal pattern classification.
To meet the demand of balancing the quality of prediction and efficiency in learning Transformer-based models, researchers proposed hierarchical structure in Duan et al. (2022), pyramidal structure in Pyraformer Liu et al. (2022a), AutoEncoders Asres et al. (2022), U-Net-like model in Yformer Madhusudhanan et al. (2023) or patches in PatchTST Nie et al. (2023).
Kernel-U-Net is a symmetric and hierarchical U-shape architecture that exponentially reduces the input length at each level, thereby concurrently decreasing the complexity involved in learning long sequences. Kernel-U-Net separates the procedure of partitioning input time series into patches from kernel manipulation, thereby providing the convenience of executing customized kernels. By replacing linear kernels with transformer or LSTM kernels, the model gains enhanced expressiveness, allowing it to capture more complex patterns and dependencies in the data. Notably, when Transformer modules are utilized in the second or higher-level layers, the computation cost remains linear, ensuring efficiency in processing.
@@ 3.1 Problem Formulation
In the context of time series forecasting, the dataset is composed of a time series of characteristics $x$ and future series $\hat{x}$. Let $x_{t}$ be the feature at the time step $t$ and $L$ the length of the look-back window. Given a historical data series $(x_{t+1},...,x_{t+L})$ of length $L$, time series forecasting task is to predict the value $(\hat{x}_{t+L+1},...,\hat{x}_{t+L+T})$ in the future T time steps. Then we can define the basic time series forecasting problem: $(\hat{x}_{t+L+1},...,\hat{x}_{t+L+T})=f(x_{t+1},...,x_{t+L})$ , where $f$ is the function that predicts the value $(\hat{x}_{t+L+1},...,\hat{x}_{t+L+T})$ in the future T time steps based on a series $(x_{t+1},...,x_{t+L})$.
@@ 3.2 Kernel U-Net
Kernel U-Net(K-U-Net) is a hierarchical and symmetric U-shape architecture. It separates the procedure of partitioning input time series into patches from kernel manipulation, thereby providing the convenience of executing customized kernels(Figure 1). More precisely, the K-U-Net encoder reshapes the input sequences into a large batch of small patches and repeatedly applies custom kernels on them until obtaining the latent vectors. Later, the K-U-Net decoder develops the batch of latent vectors into patches gradually at each layer and obtains a large batch of small patches(Figure 2). At last, K-U-Net reshapes the patches to get the final output. In the following paragraphs, we describe methods such as the hierarchical partition of the input sequence and the creation of Kernel-U-Net.
Figure 2: Hierarchical Partition
Algorithm 1 Kernel Wrapper
Input: kernel $f_{i}$, input patches $x$, input length $L_{in}$ and dimension $M_{in}$, output length $L_{out}$ , and dimension $M_{out}$
Output: Instance of Kernel Wrapper
  # Define the init function :
  def __init(inputs)__:
  # Initiation of main kernel function $f$
  $f=$ $f_{i}(L_{in},M_{in},L_{out},M_{out})$
  # Holds variables for global operation
  _unet_skip_output = None
  _unet_skip_input = None
  # Define the forward function :
  def forward(x):
  reshape $x$ into (-1, $L_{in}$,$M_{in}$)
  if _unet_skip_input is not None then
           $x=x+\_unet\_skip\_input$
  end if
  # Apply $f$ on $x$
  $y=f(x)$
  _unet_skip_output = $y$
  return $y$
@@
Let us consider an input trajectory matrix $X_{t}\in R^{L\times M}$, where $L$ is lookback window size, $M$ is feature size and $t$ is the time step. Given a list of multiples $(L_{1},...,L_{n})$ for look-back window and $(M_{1},...,M_{m})$ for feature such that $L=\prod_{1}^{n}L_{i}$ and $M=\prod_{1}^{m}M_{i}$ We reshape $X_{t}$ as small patches $X_{t,i,j}\in R^{L_{1}\times M_{1}}$ such that the concatenation of all patches is equal to its original trajectory matrix:
@@
The partitioned sequences form a large batch of patches and will be processed by kernels in the encoder. Since the operations in the decoder are symmetrical to those in the encoder, there will also be output $\hat{X}_{t+L}$ composed of generated patches $\hat{X}_{t+L,i,j}$.
@@
At the stage of the encoder, the shape of input $X$ is $(B,L,M)$. Let us note $L_{1}$ and $M_{1}$ as the length and feature size of a patch. Let us assume that $M_{hidden}$ is the unique dimension of the hidden vector at each intermediate layer to simplify the problem. We reshape $X$ into $(B,\prod_{2}^{n}L_{i},L_{1},\prod_{2}^{m}M_{i},M_{1})$ then transpose it into $(B,\prod_{2}^{m}M_{i},\prod_{2}^{n}L_{i},L_{1},M_{1})$ and reshape it into $(B*\prod_{2}^{m}M_{i}*\prod_{2}^{n}L_{i},L_{1},M_{1})$ The kernel $f_{L_{1}}$ can now process the reshaped $X$ as a large batch of patches and output a vector $Z_{L_{1}}$ in shape of $(B*\prod_{2}^{m}*M_{i}*\prod_{2}^{n}L_{i},1,M_{hidden})$. After this operation, we reshape the output $Z_{L_{1}}$ into $(B*\prod_{2}^{m}M_{i}*\prod_{3}^{n}L_{i},L_{2},M_{hidden})$ for next layer. Iteratively, the encoder of Kernel U-Net processes all the multiples in the list $(M_{m},...,M_{2},L_{n},...,L_{2})$ in reverse order, and gives the final latent vector $Z$ in the shape of $(B,1,M_{hidden})$ as a batch of the latent vector (Algorithm 2).
At the stage of the decoder, the operations are reversed. We send a latent vector Z of shape $(B,1,M_{hidden})$ to decoder $f_{M_{m}}$ and get an output $Z_{M_{m}}$ in shape of $(B,M_{m},M_{hidden})$, then we reshape it into $(B*M_{m},1,M_{hidden})$ for the next kernels. At the end of this iteration over multiples in the list $(M_{m-1},...,M_{2},L_{n},...,L_{2},L_{1})$, we have finally a vector in the shape of $(B*\prod_{2}^{m}M_{i}*\prod_{2}^{n}L_{i},L_{1},M_{1})$. The last operations are reshaping it into $(B,\prod_{2}^{m}M_{i},\prod_{2}^{n}L_{i},L_{1},M_{1})$ , transposing it into the vector in shape $(B,\prod_{2}^{n}L_{i},L_{1},\prod_{2}^{m}M_{i},M_{1})$ and reshaping it to shape $(B,L,M)$ (Algorithm 3)..